JetBlue Airways is an American low-cost airline headquartered in New York City. Its main base is John F. Kennedy International Airport. JetBlue's revenue in 2000 was $1,2401 million. By 2015, revenue had increased to $6,076.6 million

What was the geometric mean annual increase for the period?
Answer #1

Geometric Mean Rate of Increase Over Time = [(6076.60 / 1240.10) ^ (1/15)] - 1= 1.7605 - 1 = 0.7605

Rate of increase is 76.05 percent per year.